### v1

@@ Abstract.
copyright: None
@@ 1. Introduction
  • We evaluate Wukong on both our newly proposed dataset and two public benchmarks, demonstrating that it outperforms text-based safeguards and achieves comparable accuracy to image-based safeguards, while being significantly more efficient. Additionally, we test Wukong on adversarial (or jailbreaking) prompts, specifically crafted to bypass safeguards by substituting explicitly unsafe terms. The experimental results show that Wukong remains robust against such adversarial prompts. And our code is available [^1] https://anonymous.4open.science/r/Wukong-64F2.
@@ 3.1. Task Definition: Image-based NSFW Content Detection in T2I Generation
(2) $f(s,\mathcal{G},z)=\left\{\begin{aligned} &1\quad{\ if\ \mathcal{I}\ contains\ NSFW\ contents}\\ &0\quad otherwise\\ \end{aligned}\right.$
@@ 3.3. U-Net and Cross Attention
Figure 2. Visualization of Attention Maps. (a) shows an image generated with the prompt ”car” using Stable Diffusion XL. (b)–(e) display attention maps from the last cross-attention layer in U-Net’s upsample block at various denoising steps, where queries are derived from intermediate latent representations and keys from the encoded text ”car”. (f) shows an image generated with the prompt ”a naked woman”, while (g)–(j) show corresponding attention maps using the NSFW concept ”sexual” as the key. All attention maps are reshaped from 4096×14096$\times$ 14096 × 1 to $\times 64$ 64, then resized to $\times 1024$ 1024 for visualisation.
@@ 4.3. Training Objective
(14) $\mathcal{L}_{{\rm BCE}}=\sum_{i=1}^{7}\left[y_{i}\cdot\log(\hat{y_{i}})+(1-y_{i})\cdot\log(1-\hat{y_{i}})\right].$
This formulation encourages the model to learn category-specific NSFW patterns, and allows it to handle multi-label cases where multiple types of unsafe content may coexist in a single image. During inference, we apply a predefined threshold $\delta$ to the maximum value of the predicted probabilities $\max(\hat{y})$ to obtain binary decisions.
@@ 4.4. Integrated Pipeline
5:   $\hat{y}={\rm classifier}\left(\varphi\left(x_{T-t}\right)\right)$triangleright all layers in Section 4.2
@@ 5. Wukong-demons Dataset
Existing NSFW datasets for T2I generation (e.g., I2P (Schramowski et al., 2023), CoPro (Liu et al., 2024)) typically assign labels based solely on textual inputs, overlooking model-specific factors, such as how image safety may vary across different T2I models or even across random seeds for the same prompt. To the best of our knowledge, no existing dataset provides seed- and model-specific NSFW annotations based on the generated image itself.
To address this gap and support more rigorous evaluation, we introduce the Wukong-Demons dataset. Inspired by CoPro (Liu et al., 2024), it begins with seven core NSFW categories, expands to semantically related concepts (e.g., ”knife” for ”violence”), and then generates full prompts incorporating these concepts. Using GPT-4o-mini, we generate 50 visually unsafe concepts per category and 20 descriptive prompts per concept, yielding ~1,000 prompts per category. The dataset is split into training, validation, and test sets in a 7:2:1 ratio. For each prompt, we generate 10 images using different random seeds and label each image with category-specific NSFW tags using GPT-4o-mini as a vision-language model (VLM), recording both labels and seeds.
@@
This construction offers two key improvements:
  1. (1)
    We leverage a strong VLM to ensure that generated prompts describe visually unsafe content, improving alignment between prompt semantics and visual outcomes.
  2. (2)
    We assign NSFW labels based on the generated image rather than the text, enabling model- and seed-specific annotations that better reflect T2I behavior.
This two-step strategy improves visual relevance and mitigates the shortcomings of purely text-based labeling, such as overgeneralizing any mention of ”knife” as unsafe, an issue common in prior datasets (Liu et al., 2024; Li et al., 2024).
We provide a detailed description of the Wukong-Demons dataset in Appendix A, including prompt construction, data format, and statistical analysis.
6. Experiments
6.1. Datasets
6.2. Baselines
@@ 6.7. Ablation Study
Wukong
Demons
Wukongw/o Att 0.8651 0.8589
@@
(a) Wukong-Demons
(b) I2P
(c) CoPro
Figure 4. The performances comparison varying $T_{C}$
@@ 6.8.2. Efficiency
In contrast, traditional image filtering approaches require the entire diffusion process to complete before applying an image-level classifier, typically taking over 10 seconds. Thus, Wukong achieves over 5$\times$ speedup while maintaining competitive performance, highlighting its advantage for real-time or resource-constrained deployment scenarios.
@@ 7. Conclusion
In this paper, we present Wukong, a transformer-based framework for NSFW content detection in text-to-image generation. By leveraging intermediate latent representations from early denoising steps and reusing pre-trained U-Net cross-attention parameters, Wukong enables efficient and accurate classification without full image generation. Additionally, we contribute a new dataset with prompt-image pairs and category-specific NSFW labels to support future research. Extensive experiments demonstrate that Wukong significantly outperforms existing text-based filters in accuracy and offers substantial efficiency gains compared to traditional image-based approaches.
Appendix A Details of the Wukong-demons Datasets
This section provides a detailed overview of the construction process for the Wukong-demons dataset.
A.1. Textual Prompt Generation
Following prior works, we focus on seven NSFW categories: {Illegal Activity, Hate, Violence, Sexual, Self-harm, Harassment, Shocking(Schramowski et al., 2023; Liu et al., 2024; Li et al., 2024). For each category, we use GPT-4o-mini to generate 50 related concepts using the following prompt template:
Prompt Template for NSFW Concepts Generation List the related concepts about the category {category}. Namely, when describing a {category} scene, what terms or phrases are usually used? The terms must directly depict visually {category} descriptions.
For example, concepts generated for the Sexual category include: Lustful, Breasts, Naked, etc. If fewer than 50 concepts are initially produced, we regenerate additional ones using the same prompt until the total reaches or exceeds 50. All generated concepts are retained without filtering.
Next, for each concept, we generate 20 textual prompts using the following template:
To ensure compatibility with the text encoder of the Stable Diffusion pipeline (which limits inputs to 77 tokens), we restrict each prompt to under 70 words. This process ensures the generated prompts are both visually descriptive and semantically aligned with their respective NSFW categories, forming a high-quality basis for image generation and subsequent safety evaluation.
A.2. Image and Label Generation
After generating the textual prompts, we use the T2I models (e.g., Stable Diffusion XL (SDXL) model (Stability, 2023)) to generate 10 images per prompt, recording the random seed for each to ensure reproducibility. Each generated image is then labeled using GPT-4o-mini [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, which assigns category-specific safety labels along with brief justifications. The following prompt is used for labeling:
Table 6. Category-wise distribution statistics of the Wukong-demons dataset
Illegal Activity Hate Violence Sexual Self-Harm Harassment Shocking Overall
# of prompts 1020 1280 1200 1020 1040 1080 1000 7640
% of unsafe images 47.21 47.07 83.01 70.21 58.24 20.56 52.4 54.29
Table 7. NSFW image statistics across three datasets
CoPro I2P Wukong-demons
% of unsafe images 21.35 32.28 54.29
% of 1$\geq$ 1≥ 1 prompts 56.14 82.77 90.26
@@ A.2. Image and Label Generation
A.3. Statistics of Dataset
Table 7 compares the distribution of the Wukong-Demons dataset with CoPro and I2P. Here, ”of unsafe images” refers to the proportion of all generated images labeled as unsafe, while ”of 1$\geq$ 1≥ 1 prompts” indicates the proportion of prompts that generate at least one unsafe image across 10 random seeds. As shown, the Wukong-Demons dataset contains a substantially higher proportion of visually unsafe prompts compared to previous datasets.
@@ Appendix B Selection of Baseline Methods
We compare our proposed Wukong framework against seven mainstream and state-of-the-art external text-based NSFW safeguards, as introduced in Section 6.2. These baseline methods comprehensively represent the current landscape of NSFW detection approaches, spanning from simple rule-based techniques to advanced neural architectures. Specifically, they include:
  • Rule-based methods: Text Blacklist (Ronik Patel, 2024) is a keyword-matching safeguard that flags prompts containing predefined unsafe terms. This is the most lightweight and computationally efficient method. However, it is highly vulnerable to adversarial prompts, where unsafe keywords are replaced with benign alternatives to evade detection.
  • Moderation models: The OpenAI Moderation model (ope, line) and the NSFW Text Classifier (Li, [n. d.]) are well-pretrained, publicly available moderation tools designed specifically for NSFW content detection. These models are capable of capturing richer semantic information beyond simple keyword matching, and are widely adopted in both academic and industry settings.
  • Visual Language Model (VLM)-based methods: GPT-4o-mini (OpenAI, 2024) and InternVL2-8B (Chen et al., 2024) leverage the strong multimodal reasoning capabilities of recent large VLMs, which have demonstrated impressive performance in understanding both text and image inputs. Prior works (Liu et al., 2024) highlight the effectiveness of VLMs in safety-critical applications. However, these models remain susceptible to jailbreaking or adversarial prompts: prompts that have been subtly modified to bypass safeguards while retaining harmful intent, such as replacing unsafe words with syntactically benign alternatives or encoded representations (Yang et al., 2024a, c).
  • Advanced neural architectures: Latent Guard (Liu et al., 2024) and AEIOU (Wang et al., 2024) represent more recent and sophisticated NSFW detection frameworks that aim to improve robustness against adversarial inputs through specialized neural design. These approaches are specifically engineered to address the limitations of earlier moderation and VLM-based systems.
The inclusion of this diverse set of baselines ensures comprehensive coverage of current external safeguard paradigms, ranging from rule-based to semantically-aware to adversarially robust detectors.
However, we do not adopt image-based safeguards or internal safeguards as baselines in our main comparisons, as they serve fundamentally different roles or involve substantially different computational trade-offs, which we discuss in detail below.
@@ Appendix C Implementations and Settings of Baseline Methods
Visual language models (GPT-4o-mini (OpenAI, 2024) and InternVL2-8B (Chen et al., 2024)) are queried using the following prompt template:
Prompt Template for VLM-based NSFW Detection Detect whether the following user prompt contains Not Safe For Work (NSFW) content.
Prompt: {user prompt} Format: Classification: [yes — no] Reason: [Short justification]

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 9–13, 2025; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD 2026), August 9–13, 2025, Jeju Island, Republic of Koreaisbn: 979-8-4007-2258-5/2026/08doi: 10.1145/3770854.3780226ccs: Computing methodologies Computer vision tasksccs: Security and privacy Software and application security
@@ 1. Introduction
  • We evaluate Wukong on both our newly proposed dataset and two public benchmarks, demonstrating that it outperforms text-based safeguards and achieves comparable accuracy to image-based safeguards, while being significantly more efficient. Additionally, we test Wukong on adversarial (or jailbreaking) prompts, specifically crafted to bypass safeguards by substituting explicitly unsafe terms. The experimental results show that Wukong remains robust against such adversarial prompts. And our code is available [^1] https://github.com/MINGRUI001/Wukong.
@@ 3.1. Task Definition: Image-based NSFW Content Detection in T2I Generation
[…]
@@ 3.3. U-Net and Cross Attention
Figure 2. Visualization of Attention Maps. (a) shows an image generated with the prompt ”car” using Stable Diffusion XL. (b)–(e) display attention maps from the last cross-attention layer in U-Net’s upsample block at various denoising steps, where queries are derived from intermediate latent representations and keys from the encoded text ”car”. (f) shows an image generated with the prompt ”a naked woman”, while (g)–(j) show corresponding attention maps using the NSFW concept ”sexual” as the key. All attention maps are reshaped from $4096\times 1$ to $64\times 64$, then resized to $1024\times 1024$ for visualisation.
@@ 4.3. Training Objective
[…]
@@ 4.4. Integrated Pipeline
5:   $\hat{y}={\rm classifier}\left(\varphi\left(x_{T-t}\right)\right)$ $\triangleright$ all layers in Section 4.2
@@ 5. Wukong-Demons Dataset
Existing NSFW datasets for text-to-image (T2I) generation (e.g., I2P (Schramowski et al., 2023), CoPro (Liu et al., 2024)) typically assign safety labels based solely on textual prompts. This overlooks model-specific factors, as image safety can vary across different T2I models or even across random seeds for the same prompt. To the best of our knowledge, no existing dataset provides NSFW annotations grounded in the generated images themselves, with explicit consideration of model and seed variability.
To address this gap, we introduce the Wukong-Demons dataset, a large-scale benchmark with image-based, category-level NSFW annotations. Following prior work (Schramowski et al., 2023; Liu et al., 2024; Li et al., 2024), we focus on seven NSFW categories: {Illegal Activity, Hate, Violence, Sexual, Self-harm, Harassment, Shocking}. Inspired by CoPro (Liu et al., 2024), our dataset construction proceeds in three stages: concept expansion, prompt generation, and image-level labeling.
Concept Expansion.
We use a similar concept expansion method as in Latent guard (Liu et al., 2024), where the prompt template is shown in Appendix A.1.
If fewer than 50 concepts are produced, we regenerate additional ones until the target number is reached. All generated concepts are retained without manual filtering.
Prompt Generation.
For each concept, we generate 20 descriptive prompts using the following template:
We limit each prompt to fewer than 70 words to ensure compatibility with the Stable Diffusion text encoder (77-token limit). This results in approximately 1,000 prompts per category. The dataset is split into training, validation, and test sets in a 7:2:1 ratio.
Image Generation and Labeling.
For each prompt, we generate 10 images using a T2I model (e.g., SDXL (Stability, 2023)), varying random seeds and recording them for reproducibility. Each image is labeled with category-specific NSFW tags using GPT-4o-mini as a vision–language model (VLM), based on the following prompt:
To improve label reliability, we additionally use OpenAI’s multimodal moderation model (ope, line) as auxiliary supervision. An image is marked unsafe if the highest predicted unsafe-category probability exceeds 0.5. When the moderation output conflicts with the VLM label, we convert the moderation result into text and provide it to the VLM for reconsideration. We further manually verify approximately 10% of generated images—primarily those with conflicting labels—and observe high agreement between the VLM’s final decisions and human judgments.
Dataset Advantages.
This construction offers two key benefits:
  1. (1)
    Using a strong VLM ensures that prompts correspond to visually unsafe content, improving alignment between textual semantics and generated images.
  2. (2)
    Assigning labels based on generated images—rather than prompts alone—enables model- and seed-specific annotations that more accurately reflect T2I behavior.
This image-grounded labeling mitigates common issues in text-only datasets, such as overgeneralizing any mention of terms like ”knife” as unsafe (Liu et al., 2024; Li et al., 2024).
Table 1 compares the distribution of the Wukong-Demons dataset with CoPro and I2P. Here, ”% of unsafe images” refers to the proportion of all generated images labeled as unsafe, while ”% of $\geq 1$ prompts” indicates the proportion of prompts that generate at least one unsafe image across 10 random seeds. As shown, the Wukong-Demons dataset contains a substantially higher proportion of visually unsafe prompts compared to previous datasets.
Table 1. NSFW image statistics across three datasets
CoPro I2P Wukong-demons
% of unsafe images 21.35 32.28 54.29
% of $\geq 1$ prompts 56.14 82.77 90.26
Additional examples are provided in Appendix A.2.
6. Experiments
6.1. Datasets
6.2. Baselines
@@ 6.6. Performance on Adversary Prompts
6.7. Per-category Performance
Table 4. Top-1 category performance in terms of ROC AUC and accuracy
Metric Wukong-demons I2P CoPro
ROC AUC 0.9429 0.8578 0.9211
Accuracy 0.9386 0.8297 0.9175
We report per-category top-1 performance in Table 4, where a prediction is considered correct if the highest-probability predicted label matches one of the ground-truth unsafe categories. The resulting accuracies, 0.9386 on Wukong-demons, 0.8297 on I2P, and 0.9175 on CoPro, are close to the corresponding overall accuracies (0.9452, 0.8582, and 0.9231, respectively), indicating that the classifier maintains consistent performance at the category level.
@@ 6.8. Ablation Study
Wukong - demons Wukongw/o Att 0.8651 0.8589
@@ 6.9.2. Efficiency
[…]
6.10. Impact of $\delta$
We study the effect of the threshold $\delta$ on NSFW detection performance. Since ROC AUC is threshold-independent, we report F1-score on all three datasets.
Figure 5. F1-score under different values of $\delta$.
As shown in Figure 5, performance peaks around $\delta\in[0.5,0.6]$ and remains stable over a wide range. For instance, on Wukong-demons, the F1-score stays above 0.88 when $\delta\in[0.4,0.7]$, demonstrating low sensitivity to $\delta$ and robust performance in practice.
@@ 7. Conclusion
In this paper, we present Wukong, a transformer-based framework for NSFW content detection in text-to-image generation. By leveraging intermediate latent representations from early denoising steps and reusing pre-trained cross-attention modules within the diffusion model, Wukong enables efficient and accurate classification without requiring full image generation. We also introduce a new dataset of prompt–image pairs with category-level NSFW annotations to facilitate future research. Extensive experiments demonstrate that Wukong significantly outperforms existing text-based filters in accuracy while achieving substantial efficiency gains compared to traditional image-based approaches. Finally, recent diffusion models such as DiT (Peebles and Xie, 2023) adopt fully transformer-based backbones instead of the U-Net-style architectures used in Stable Diffusion. While this work focuses on U-Net–based diffusion models with cross-attention mechanisms, extending Wukong to DiT-style architectures is a promising direction for future work.
Acknowledgements.
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund (Tier 2 Award MOE-T2EP20224-0011 and Tier 1 Award (RG20/24)). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendix A Details of the Wukong-demons Dataset
This section provides some supplementary details about Wukong-demons dataset.
A.1. Concept Expansion Prompt
We use GPT-4o-mini to generate 50 related concepts using the following prompt template:
Prompt Template for NSFW Concepts Expansion List the related concepts about the category {category}. Namely, when describing a {category} scene, what terms or phrases are usually used? The terms must directly depict visually {category} descriptions.
A.2. Examples of Wukong-dataset
@@ Appendix B Selection of Baseline Methods
We compare our proposed Wukong framework against seven mainstream and state-of-the-art external text-based NSFW safeguards, as introduced in Section 6.2. These baseline methods comprehensively represent the current landscape of NSFW detection approaches, spanning from simple rule-based techniques to advanced neural architectures. However, we do not adopt image-based safeguards or internal safeguards as baselines in our main comparisons, as they serve fundamentally different roles or involve substantially different computational trade-offs, which we discuss in detail below.
@@ Appendix C Implementations and Settings of Baseline Methods
Visual language models (GPT-4o-mini (OpenAI, 2024) and InternVL2-8B (Chen et al., 2024)) are queried using the following prompt template shwon in our repository [^2] https://github.com/MINGRUI001/Wukong.